# QF-tuner: Breaking Tradition in Reinforcement Learning

### v3

@@ Abstract
Hyperparameter tuning in reinforcement learning algorithms refers to choosing the optimal parameters that may increase the algorithm’s performance. Manual or random hyperparameter tuning methods can be problematic, as even slight variations in their values can result in significantly different outcomes in the learning process. In this paper, we propose a new method, QF-tuner, for automatic hyperparameter tuning in the Q-learning algorithm using the FOX optimization algorithm (FOX). A new objective function has been proposed for the FOX, prioritizing reward over learning error and time. QF-tuner starts by running the FOX and tries to minimize the fitness value derived from observations at each iteration by executing the Q-learning algorithm. The proposed method has been evaluated using two control tasks from the OpenAI Gym: CartPole and FrozenLake. The empirical results of the QF-tuner on the CartPole control task show a reward of 499, and on the FrozenLake control task, a reward of 1. These results indicate that the QF-tuner outperforms other optimization algorithms. On the FrozenLake control task, there was a 36% increase in reward with a 26% reduction in learning time; on the CartPole control task, there was a 57% increase in reward with a 20% decrease in learning time. Thus, the QF-tuner is an essential method for hyperparameter tuning in reinforcement learning algorithms, enabling more effective solutions to control task problems.
@@ 1 Introduction
Reinforcement learning (RL) involves selecting the actions that achieve the maximum rewards in each state of the problem’s environment. This behavior is learned through interactions with the environment by observing its trial-and-error responses. It is similar to how children explore their surroundings and discover the actions that lead to goal achievement [1]. Q-learning is considered a standard algorithm in RL, which is model-free off-policy, widely used, and simple to implement to solve finite markov decision processess (MDPs). RL algorithms solve problems, such as gaming, controlling, and scheduling, by moving from one state to another based on an approximating action-value function in the problem’s environment to select the optimal action that make a correct decision in the current state [2, 3]. The process of approximating the action-value function is repeated several episodes to increase the reward and find the optimal policy [4]. The FOX optimization algorithm (FOX) is a recently proposed metaheuristic optimization algorithm that takes inspiration from nature by mimicking how foxes hunt prey [5]. However, in RL algorithms, hyperparameters are set before learning that controls the learning process. They affect how the algorithm explores and learns from the environment, such as the learning rate, discount factor, and exploration strategy [6, 7]. While RL algorithms can yield maximum rewards; they require overhead tuning the hyperparameter to achieve optimum performance [8]. The difficulty is in determining the most effective values for each hyperparameter, as even minor adjustments might influence the learning process in any RL algorithm. However, manual hyperparameter tuning can be complex and time-intensive. Several methods have been suggested for hyperparameter tuning manually, randomly, or semi-automatic adjusting values for each parameter [9]. Nevertheless, these methods may not generally apply to all RL problems due to environmental variables and conditions variations.
@@
This paper employed the FOX for automatic hyperparameter tuning in the Q-learning algorithm to find the optimal parameter that results in maximum rewards and the lowest learning time (steps). This method, QF-tuner, solves the problems associated with manual hyperparameter tuning in RL algorithms. It offers a robust and effective solution that addresses the challenges of hyperparameter tuning. Furthermore, a new objective function has been proposed and utilized in the FOX. it includes reward, error, and learning time and prioritizes reward over error and learning time. However, the contributions of this paper are as follows:
@@ 2 Related Work
In the past few years, many investigations have focused on solving control tasks using RL, including the conventional or improved Q-learning algorithm. The authors of [11] proposed a framework for reducing the burden of hyperparameter tuning by using random feature approximation to efficiently compute state features and a factorial policy that allows fast processing of many discrete actions. The proposed framework was able to reduce the overhead of hyperparameter tuning in the context of large-scale control tasks. Furthermore, the authors [12] suggested two Q-learning-based methods for solving the CartPole control task. They used mean squared error (MSE) and mean absolute error (MAE) as reward objective functions. The evaluation focused on how well each method balanced the pole, intending to maximize the reward. Their results show that Q-learning with MSE tends to ignore outliers while with MSE gives more weight to outlier values. Moreover, the authors [13] designed a software package with a scalable GA to explore the best hyperparameters in deep RL. This package has been tested on OpenAI Gym environments and compared to the Bayesian approach. The study’s conclusions show that this approach is effective at maximizing reward with fewer episodes
The authors [14] proposed a hybrid model combining the bees algorithm (BA) with Q-learning to improve feature selection in coronary heart disease diagnosis. The proposed method was experimented on multiple datasets and showed improved accuracy compared to conventional algorithms. Another study by [15] presented a comparison study on the FrozenLake control task between the Q-learning algorithm and cognitive instance-based learning (IBL) algorithms. The Q-learning and  IBL agents were tested on different grid sizes of Frozen Lake. The IBL algorithm surpassed the Q-learning algorithm by demanding less learning time and better generalizing. However, the study did not address Q-learning’s limitations, underscoring the challenge of developing an enhanced version capable of effectively solving FrozenLake and other control tasks. Moreover, the authors [16] suggested Convergence-AwarE SAmpling with scReening (CAESAR) aggregation scheme. CAESAR blends convergence-aware sampling with a screening mechanism to integrate knowledge from agents learning in similar MDPs, improving overall learning efficiency. The results of the Frozen Lake control task validate CAESAR’s significance in improving agent learning across various control task environments.
@@
However, existing hyperparameter tuning methods like particle swarm optimization (PSO)genetic algorithms (GA), and BA often struggle to escape local optima in complex control tasks [14, 17, 18]. These methods rely on fixed or semi-adaptive strategies, limiting their effectiveness. In contrast, the QF-tuner introduces an automatic tuning method that adjusts hyperparameters based on novel objective functions, overcoming these limitations and improving stability and efficiency, especially in tasks where other algorithms face convergence issues.
@@ 3 Materials and Methods
This section presents materials (control tasks) and the proposed method (QF-tuner). It starts by explaining the OpenAI Gym environment control tasks—FrozenLake and CartPole. Subsequent sections will provide a detailed background of the Q-learning algorithm, FOX, and the QF-tuner.
@@ 3.1 Materials
This paper used two control tasks—FrozenLake and CartPole—developed by OpenAI Gym to evaluate the QF-tuner. In the FrozenLake control task, an agent has to cross a grid world over a FrozenLake. This task introduces obstacles, creating a challenging environment [19]. Furthermore, in the CartPole environment, an inverted pendulum on a moving cart needs to be stabilized. This task is common when evaluating the ability of RL algorithms to maintain stability in a continuous state space [20]. However, QF-tuner uses these tasks to guarantee a standardized evaluation framework for evaluating the performance. Moreover, this paper advances experiential and makes relevant comparisons with previous research findings using other methods on these control task environments.
@@ 3.2 Q-learning Algorithm
In the learning process, the Q-learning’s agent obtains the reward $R$ after taking the action $A$ in the current state $S$, then moves to the next state $S_{t+1}$. Additionally, the Q-learning algorithm tries to find the best course of actions that follow policy in the MDP by determining the optimal function of action-value using the update rule from eqs. 1 and 2 [25, 26, 27]. Moreover, the agent interacts with the environment to observe the reward and updates the action-value function to improve its policy estimates over time. Agent interactions process iteratively, allowing the algorithm to modify its policy based on observations of the problem’s environment. Finally, the agent should have the ability to select the best action (optimal policy) in a given state using the estimated Q-table by the following eq. 3 [28]:
@@ 3.3 FOX optimization algorithm
FOX is a new optimization algorithm inspired by red foxes’ hunting behavior that aims to search for the the best solution with the best fitness value [5]. FOX has many search agents working together iteratively, each trying to find the best fitness value during the search. It involves two major phases: exploration and exploitation. In exploration, the Fox agent uses a random walk strategy to locate and catch prey, utilizing its ability to detect ultrasound. During its search for prey, the Fox agent may listen for sounds indicating where its prey is. In exploitation, the Fox agent can hear ultrasound, so it takes time for the sound of its prey to reach it. The speed of sound is 343 meters per second [30]. However, since this speed is always the same, FOX uses a different method to measure it. The Fox agent jumps when it thinks it can catch its prey based on how long the sound takes to reach it. Thus, the Fox agent’s ability to catch prey depends on the time it takes for the sound to reach them while jumping [31]. The following equations described the primary steps of FOX in detail.
@@
The variable $DSF_{i}$ represents the sound distance from the Fox agent. The sound speed in the medium is represented by the variable $S$ and calculated by eq. 4. The variable $T$ represents an arbitrary number in the $[0,1]$ range. The variable $i$ represents the current iteration of the optimization process [32]. Moreover, The distance of a Fox agent from its prey is defined as follows [33]:
@@
Once a Fox agent has determined the distance between itself and its prey, it should jump and catch the prey accurately. The Fox agent must calculate the required jump height using the eq. 7 to perform this jump [5, 34].
@@
where the value of $c_{1}$ and $c_{2}$ are 0.180 and 0.820, respectively. These values have been determined based on the Fox agent’s jump. The direction of the jump is randomly chosen to be either northeast or opposite direction. However, the Fox agent continuously moves to a new position in exploration and exploitation. In exploitation, the new position is determined using either eq. 8 or eq. 9.
@@
$X_{i+1}=\text{Best}X\times\text{rnd}(1,\text{dim})\times\min(tt)\times a$ (10)
The variable $tt$ is the time average, which is equal to the summation of the variable $T$ used in eq. 4 divided by the dimension of the problem (dim). The exploration is controlled by variable $a$ that is computed by $2\times[i-(i/max(iter))]$. However, the source code of FOX implementation can be accessed by https://github.com/Hardi-Mohammed/FOX. The next subsequent section will present an explanation of the QF-tuner.
@@ 3.4 Hyperparameter Tuning Method (QF-tuner)
The Q-learning algorithm has several hyperparameters that affect the agent’s performance and behavior regarding the learned policy. From eqs. 2 and 1, the following hyperparameters can be noticed: step size ($\alpha$) and discount factor ($\gamma$). They should be set before the beginning of the learning process. The $\alpha$ is crucial in updating the action-values function at each time step. It determines the extent to which new information overrides old information, and its value is typically within the range $[0,1]$. A lower $\alpha$ value means that the algorithm will be less sensitive to new information and will take longer to converge, while a higher value results in faster learning but may lead to instability. Therefore, choosing the correct value of $\alpha$ is essential for achieving the best performance in any learning algorithm [35]. In comparison, the $\gamma$ affects future reward for an agent’s decisions and its value range $[0,1]$. A more significant value (close to 1) indicates that the agent emphasizes future rewards, while a smaller value means that the agent gives less importance to future rewards. This parameter controls the balance between immediate and future rewards, affecting the agent’s decision-making process [36].
@@
$\textit{fitness}=\sum_{i=1}^{n}(2R-e)\times\frac{1}{st}$ (11)
where the $e$ variable represents the temporal difference error computed by the eq. 1 and the value of $st$ represents the amount of time (learning steps) the Fox agent takes to reach the best solution. Indeed, this equation is calculated on the last quarter of episodes from the entire episodes to ensure the stability of the learning process, so $i=episodes-episodes/4$. For example, if the number of episodes = 200, the fitness calculation starts from the episode 150.
The QF-tuner employed the FOX with Q-learning algorithms in a new model that automates the tuning of hyperparameters without human intervention. Setting up the QF-tuner involves determining the number of iterations $max\_iter$, the number of Fox agents $g$, and the number of runs $n$, as illustrated in fig. 1. The FOX is employed, and its initial solution serves as the initial hyperparameters ($\alpha$ and $\gamma$) of the Q-learning algorithm. Next, The Q-learning algorithm starts learning iteratively based on the number of episodes, during which reward, learning error, and convergence time values are computed to derive the fitness value using eq. 11. The Fox agent’s solution is updated if $max\_iter$ or appropriate convergence is not reached. Otherwise, the QF-tuner starts again until the number of runs ($n$) is achieved to prevent getting stuck in the local optima. Finally, the optimal solution consists of the best set of hyperparameters ($\alpha$ and $\gamma$), providing robust and practical solutions for real-world problem-solving, without the need for manual or random tuning.

### v4

@@ Abstract
In reinforcement learning algorithms, the hyperparameters tuning method refers to choosing the optimal parameters that may increase the overall performance. Manual or random hyperparameter tuning methods can lead to different results in the reinforcement learning algorithms. In this paper, we propose a new method called QF-tuner for automatic hyperparameter tuning in the Q-learning algorithm using the FOX optimization algorithm (FOX). Furthermore, a new objective function has been employed within FOX that prioritizes reward over learning error and time. QF-tuner starts by running the FOX and tries to minimize the fitness value derived from observations at each iteration by executing the Q-learning algorithm. The proposed method has been evaluated using two control tasks from the OpenAI Gym: CartPole and FrozenLake. The empirical results indicate that the QF-tuner outperforms other optimization algorithms, such as particle swarm optimization (PSO)bees algorithm (BA)genetic algorithms (GA), and the random method. However, on the FrozenLake task, the QF-tuner increased rewards by 36% and reduced learning time by 26%, while on the CartPole task, it increased rewards by 57% and reduced learning time by 20%. Thus, the QF-tuner is an essential method for hyperparameter tuning in Q-learning algorithms, enabling more effective solutions to control task problems.
@@ 1 Introduction
Reinforcement learning (RL) is a branch of machine learning in which the agent learns how to make a perfect decisions by interacting with an environment. It aims to identify optimal actions based on the results of these interactions to develop strategies that maximize long-term rewards. Moreover, RL involves selecting the actions that achieve the maximum rewards in each state of the problem’s environment. This behavior is learned through interactions with the environment by observing its trial-and-error responses. It is similar to how children explore their surroundings and learn from experiences to develop effective strategies through repeated practice [1, 2].
markov decision processes (MDP) is formally defined as a tuple $(S,A,P,R,\gamma)$. The $S$ represents a finite set of states, and $A$ denotes a finite set of actions available to an agent. When the agent selects an action $a\in A$ while in a current state $s\in S$, the probability of transitioning to the subsequent state $s^{\prime}\in S$ is governed by the state-transition probability function $P:S\times A\times S\rightarrow[0,1]$. Upon transitioning from state $s$ to state $s^{\prime}$ after performing action $a$, the agent receives an immediate numerical reward defined by the reward function $R:S\times A\times S\rightarrow\mathbb{R}$. The discount factor $\gamma\in[0,1)$ indicates how much the agent values immediate rewards in relation to future rewards. Thus, reflecting the importance assigned to long-term outcomes through the decision-making process [3]. Q-learning is considered a common algorithm in RL, which is model-free, off-policy, widely used, and simple to implement used to solve finite MDPs. It solves problems by transitioning between states based on an approximate action-value function. This approach allows the Q-learning to select actions that maximize cumulative rewards within the problem environment [4]. However, Q-learning is utilized in many domains, including gaming, control systems, robotics, and task scheduling [5, 6].
@@
The FOX optimization algorithm (FOX) is a recent optimization algorithm proposed by Mohammed and Rashid in 2022. FOX inspired by the behaviors of red foxes, such as walking, jumping, and hunting for prey. This algorithm mathematically simulates these natural behaviors. It works as a population-based, where multiple independent agents, known as Fox agents (FAs), explore potential solutions. FOX has shown superior performance compared to established optimization algorithms like PSO, GA, GWO, FDO, and WOA [7]. However, in RL algorithms, hyperparameters are set before the beginning of learning process. They used to control the learning behavior of the RL algorithm. They affect how the algorithm explores and exploit within the environment, such as the learning rate and discount factor [8, 9]. While RL algorithms can yield maximum rewards; they require overhead tuning the hyperparameter to achieve optimum performance [10]. The difficulty is in determining the most effective values for each hyperparameter, as even minor adjustments might influence the learning process in any RL algorithm. However, manual or random hyperparameter tuning can be complex and time-intensive process. Several methods have been suggested for hyperparameter tuning manually, randomly, or semi-automatic adjusting values for each parameter [11]. Furthermore, these methods may not generally apply to all RL problems due to environmental variables and conditions variations.
@@
This paper employed the FOX for automatic hyperparameter tuning in the Q-learning algorithm to find the optimal hyperparameters that results in maximum rewards and the lowest learning time (steps). This method, QF-tuner, solves the problems associated with manual hyperparameter tuning in RL algorithms. It offers a robust and effective solution that addresses the challenges of hyperparameter tuning. Furthermore, a new objective function has been proposed and utilized in the FOX. it includes reward, error, and learning time and prioritizes reward over error and learning time. However, the contributions of this paper are as follows:
@@ 2 Related Work
In recent years, many investigations have focused on solving control tasks using RL, particularly with Q-learning and its variants. While Q-learning is a powerful algorithm for such tasks, its performance heavily relies on hyperparameter tuning, which can be a time-consuming and challenging process. Many existing approaches have attempted to reduce this burden, but they still face limitations in large-scale or complex control environments.
For instance, the authors of [13] proposed a framework to reduce the burden of hyperparameter tuning by using random feature approximation for efficient state feature computation and a factorial policy for fast processing of discrete actions. Although their approach effectively reduced overhead for large-scale tasks, it does not directly address the challenge of automatically optimizing hyperparameters in Q-learning, which is still a bottleneck for performance improvement. Furthermore, the study by [14] suggested two Q-learning-based methods for the CartPole task. They used mean squared error (MSE) and mean absolute error (MAE) as reward objective functions. While their results demonstrated that Q-learning with MSE tends to ignore outliers, and MAE gives more weight to outliers, these methods still require manual adjustment of hyperparameters for optimal performance, which can be inefficient to suboptimal solutions. Further advancements were made by the authors of [15], who proposed a scalable GA for hyperparameter exploration in deep RL. However, while this method was effective in maximizing rewards in fewer episodes, it still faced challenges in terms of computational cost and the ability to escape local optima in complex tasks. Additionally, some hybrid approaches, such as the combination of the BA and Q-learning in [16], have been proposed to enhance feature selection in classification tasks, such as coronary heart disease diagnosis. However, while these hybrid models showed improved accuracy in specific domains, they still rely on fixed or semi-adaptive optimization strategies that struggle with the non-stationary nature of RL tasks.
In the case of the FrozenLake task, [17] compared Q-learning to cognitive IBL algorithms and found that IBL outperformed Q-learning in terms of reduced learning time and better generalization. However, this study highlighted the challenge of developing more effective Q-learning-based methods that can handle such tasks without facing the same limitations in time efficiency and generalization. Moreover, the CAESAR aggregation scheme [18] aimed to improve learning efficiency by blending convergence-aware sampling with a screening mechanism. Although it showed positive results in enhancing learning across different environments, it did not address the underlying issue of efficiently tuning the hyperparameters in Q-learning itself, which remains a key challenge for achieving optimal performance in control tasks.
@@
Despite these advances, conventional hyperparameter tuning methods like PSOGA, and BA often struggle to escape local optima, especially in the context of complex RL tasks [16, 19, 20]. These methods rely on fixed or semi-adaptive strategies, which limit their effectiveness, particularly in dynamic or highly variable task environments. In contrast, the QF-tuner introduces a novel, automatic hyperparameter tuning approach for Q-learning that leverages the FOX. This method overcomes the limitations of existing techniques by dynamically adjusting hyperparameters based on an objective function designed to prioritize reward maximization while minimizing learning error and time. By addressing the specific challenges of hyperparameter tuning in Q-learning, QF-tuner provides a more stable and efficient solution, particularly in tasks where other optimization algorithms struggle with convergence.
@@ 3 Materials and Methods
This section presents materials (control tasks) and the proposed method (QF-tuner). It starts by explaining the OpenAI Gym environment control tasks—FrozenLake and CartPole. Subsequent sections will provide a detailed background of the Q-learning algorithm, FOX, and the QF-tuner. However, the notation list of the symbols used is provided in Table 1.
Table 1: Notation List
Symbol Description
$\delta_{t}$ Temporal difference error
$R_{t+1}$ Reward at time step $t+1$
$\gamma$ Discount factor
$\alpha$ Step size
$S_{t}$ Current state
$S_{t+1}$ Next state
$A_{t}$ Current action
$a$ Selected action
$Q(S_{t},A_{t})$ Action-value function for state $S_{t}$ and action $A_{t}$
$Q^{*}(S_{t},A_{t})$ Optimal action-value function for state $S_{t}$ and action $A_{t}$
$\pi^{*}$ Optimal policy
$T_{i}$ Arbitrary number in the range [0, 1] for iteration $i$
$S$ Sound speed
$DSF_{i}$ Sound distance from the FA
$DFP_{i}$ Distance from prey
$Jump_{i}$ Jump height for the FA
$X_{i+1}$ Position of the FA at iteration $i+1$
$c_{1}$ Constant value for exploitation step 1
$c_{2}$ Constant value for exploitation step 2
$tt$ Time average
$ca$ Exploration control factor
dim Dimension of the problem
$st$ Learning steps or time taken to reach the best solution
$n$ Number of runs for the tuning process
$g$ Number of FA s
$max\_iter$ Maximum number of iterations for optimization
@@ 3.1 Materials
This paper used two control tasks—FrozenLake and CartPole—developed by OpenAI Gym to evaluate the QF-tuner. In the FrozenLake control task, an agent has to cross a grid world over a FrozenLake. This task introduces obstacles, creating a challenging environment [21]. Furthermore, in the CartPole environment, an inverted pendulum on a moving cart needs to be stabilized. This task is common when evaluating the ability of RL algorithms to maintain stability in a continuous state space [22]. However, QF-tuner uses these tasks to guarantee a standardized evaluation framework for evaluating the performance. Moreover, this paper advances experimental and makes relevant comparisons with previous research findings using other methods on these control task environments.
@@ 3.2 Q-learning Algorithm
In the learning process, the Q-learning’s agent obtains the reward $R$ after taking the action $A$ in the current state $S$, then moves to the next state $S_{t+1}$. Additionally, the Q-learning algorithm tries to find the best course of actions that follow policy in the MDP by determining the optimal function of action-value using the update rule from Equations 1 and 2 [25, 26, 27]. Moreover, the agent interacts with the environment to observe the reward and updates the action-value function to improve its policy estimates over time. Agent interactions process iteratively, allowing the algorithm to modify its policy based on observations of the problem’s environment. Finally, the agent should have the ability to select the best action (optimal policy) in a given state using the estimated Q-table by the following Equation 3 [28]:
@@ 3.3 FOX optimization algorithm
FOX is a new optimization algorithm inspired by red foxes’ hunting behavior that aims to search for the the best solution with the best fitness value [7]. FOX has many search agents working together iteratively, each trying to find the best fitness value during the search. It involves two major phases: exploration and exploitation. In exploration, the FA uses a random walk strategy to locate and catch prey, utilizing its ability to detect ultrasound. During its search for prey, the FA may listen for sounds indicating where its prey is. In exploitation, the FA can hear ultrasound, so it takes time for the sound of its prey to reach it. The speed of sound is 343 meters per second [30]. However, since this speed is always the same, FOX uses a different method to measure it. The FA jumps when it thinks it can catch its prey based on how long the sound takes to reach it. Thus, the FA’s ability to catch prey depends on the time it takes for the sound to reach them while jumping [31]. The following equations described the primary steps of FOX in detail.
@@
The variable $DSF_{i}$ represents the sound distance from the FA. The sound speed in the medium is represented by the variable $S$ and calculated by Equation 4. The variable $T$ represents an arbitrary number in the $[0,1]$ range. The variable $i$ represents the current iteration of the optimization process [32]. Moreover, The distance of a FA from its prey is defined as follows [33]:
@@
Once a FA has determined the distance between itself and its prey, it should jump and catch the prey accurately. The FA must calculate the required jump height using the Equation 7 to perform this jump [7, 34].
@@
where the value of $c_{1}$ and $c_{2}$ are 0.180 and 0.820, respectively. These values have been determined based on the FA’s jump. The direction of the jump is randomly chosen to be either northeast or opposite direction. However, the FA continuously moves to a new position in exploration and exploitation. In exploitation, the new position is determined using either Equation 8 or Equation 9.
@@
$X_{i+1}=\text{Best}X\times\text{rnd}(1,\text{dim})\times\min(tt)\times ca$ (10)
The variable $tt$ is the time average, which is equal to the summation of the variable $T$ used in Equation 4 divided by the dimension of the problem (dim). The exploration is controlled by variable $ca$ that is computed by $2\times[i-(i/max(iter))]$. However, the source code of FOX implementation can be accessed by https://github.com/Hardi-Mohammed/FOX. The next subsequent section will present an explanation of the QF-tuner.
@@ 3.4 Hyperparameter Tuning Method (QF-tuner)
The Q-learning algorithm has several hyperparameters that affect the agent’s performance and behavior regarding the learned policy. From Equations 2 and 1, the following hyperparameters can be noticed: step size ($\alpha$) and discount factor ($\gamma$). They should be set before the beginning of the learning process. The $\alpha$ is crucial in updating the action-values function at each time step. It determines the extent to which new information overrides old information, and its value is typically within the range $[0,1]$. A lower $\alpha$ value means that the algorithm will be less sensitive to new information and will take longer to converge, while a higher value results in faster learning but may lead to instability. Therefore, choosing the correct value of $\alpha$ is essential for achieving the best performance in any learning algorithm [35]. In comparison, the $\gamma$ affects future reward for an agent’s decisions and its value range $[0,1]$. A more significant value (close to 1) indicates that the agent emphasizes future rewards, while a smaller value means that the agent gives less importance to future rewards. This parameter controls the balance between immediate and future rewards, affecting the agent’s decision-making process [36].
@@
$\textit{fitness}=\sum_{i=1}^{n}(2R-\delta)\times\frac{1}{st}$ (11)
where $\delta$ represents the temporal difference error computed by the Equation 1 and the value of $st$ represents the amount of time (learning steps) the FA takes to reach the best solution. Indeed, this equation is calculated on the last quarter of episodes from the entire episodes to ensure the stability of the learning process, so $i=episodes-episodes/4$. For example, if the number of episodes = 200, the fitness calculation starts from the episode 150.
The QF-tuner employed the FOX with Q-learning algorithms in a new model that automates the tuning of hyperparameters without human intervention. Setting up the QF-tuner involves determining the number of iterations $max\_iter$, the number of FA s $g$, and the number of runs $n$, as illustrated in Figure 1. The FOX is employed, and its initial solution serves as the initial hyperparameters ($\alpha$ and $\gamma$) of the Q-learning algorithm. Next, The Q-learning algorithm starts learning iteratively based on the number of episodes, during which reward, learning error, and convergence time values are computed to derive the fitness value using Equation 11. The FA’s solution is updated if $max\_iter$ or appropriate convergence is not reached. Otherwise, the QF-tuner starts again until the number of runs ($n$) is achieved to prevent getting stuck in the local optima. Finally, the optimal solution consists of the best set of hyperparameters ($\alpha$ and $\gamma$), providing robust and practical solutions for real-world problem-solving, without the need for manual or random tuning.